Suppose you are buying your first home for $330,000, and you have $15,000 for your down payment. You have arranged to finance the remainder with a 30-year, monthly payment, amortized mortgage at a 6.5% nominal interest rate, with the first payment due in one month. What will your monthly payments be?

Select the correct answer.

a. $1,994.21
b. $1,987.81
c. $1,991.01
d. $1,997.41
e. $2,000.61
Answer #1

Remainder amount = 330,000 - 15,000 = 315,000

Rate = 6.5% / 12 = 0.541667%

Number of periods = 30 * 12 = 360

Present value = Monthly payments = [ 1 - 1 / ( 1 + r)n] / r

315,000 = Monthly payments * [ 1 - 1 / ( 1 + 0.00541667)360] / 0.00541667

315,000 = Monthly payments * 158.210754

Monthly payments = $1,991.01

Keys to use in a financial calculator:

2nd I/Y 12

I/Y 6.5

N 360

PV 315,000

CPT PMT